OpenStudy (anonymous):

Solve the equation to find the value of x. 6 - 6x = 5x - 9x - 2

OpenStudy (haseeb96):

use again same thing as i have told u in second last question

OpenStudy (haseeb96):

and now show me the work

OpenStudy (anonymous):

This one is bigger than the other one though o.o

OpenStudy (anonymous):

>.<

OpenStudy (haseeb96):

6 +2 = -4x +6x 8 = 2x 8/2 =x x = ??

OpenStudy (anonymous):

X = 4?

OpenStudy (haseeb96):

yes it is correct

OpenStudy (haseeb96):

thanks